OpenStudy (anonymous):

simplify x+6/12 over x-8/10

OpenStudy (anonymous):

you mean.....(x+6)(x+8)/30?

OpenStudy (anonymous):

\[\frac{(x+6)10}{12(x -8)}\]

OpenStudy (anonymous):

im so confuesd?

OpenStudy (anonymous):

\[\frac{\frac{x+6}{12}}{\frac{x-8}{10}}\] this is the question right ?>

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if it is then simplification i've done is correct

OpenStudy (anonymous):

its not one of my choices?

OpenStudy (anonymous):

ok I'm giving you much simplified \[\frac{(x +6)5}{(x -8)6}\]

OpenStudy (anonymous):

thanks thats it!

OpenStudy (anonymous):

thanks!